Evaluate the following integral by first converting to an integral in polar coordinates.

X^2dxdy...where x varies from -2 to 0 and y varies from -√(4-y^2) to √(4-y^2)


1
Expert's answer
2021-05-07T09:09:41-0400

Given, the integral "\int_{-2} ^{0} \int_{-\sqrt{4-y^2}} ^{\sqrt{4-y^2}} x^2 dxdy" .

From the integral limit, the curves are

"y=0" , ..........................................1

"y=-2", ....................................2

"x^2+y^2=4" . ..............................3

Convert cartesian to polar coordinate,

"x=rcos\theta," ................................4

"y=rsin\theta," .................................5

"r=\sqrt{x^2+y^2}", ........................6

"\theta=tan^{-1}\frac{y}{x}" ..............................7

From equation 3 and 6,

"r=2"

Therefore, "0\le r \le 2 \ \space and \space 0\le\theta\le2\pi".

Then, the integral

"\int_{-2} ^{0} \int_{-\sqrt{4-y^2}} ^{\sqrt{4-y^2}} x^2 dxdy \newline =\int_{0} ^{2\pi} \int_{0} ^{2} r^2cos^2\theta(rdr d\theta)\newline=\int_{0} ^{2\pi} \int_{0} ^{2} r^3cos^2\theta dr d\theta\newline
=\int_{0} ^{2\pi} [\frac{r^4}{4}]_{0}^{2}\space cos^2\theta dr d\theta\newline
=\int_{0} ^{2\pi} 1. cos^2\theta d\theta\newline
=\int_{0} ^{2\pi} \frac{1+cos2\theta}{2} d\theta\newline
=\frac{1}{2}[\theta+\frac{sin2\theta}{2}]_{0}^{2\pi}\newline
=\pi"

Thus, the value of the given integral is "\pi".